Given the probability distribution below, calculate the expected rate of return for stock A and stock...

  1. Given the probability distribution below, calculate the expected rate of return for stock A and stock B

Rate of return (%)

Probability

Stock A

Stock B

0.1

10

35

0.2

2

0

0.4

12

20

0.2

20

25

0.1

38

45

  1. Stock A = 14%; Stock B = 21%
  2. Stock A = 23% ; Stock B = 12%
  3. Stock A = 25%; Stock B  = 15%
  4. Stock A =31% ; Stock B = 27%
Answer #1

Expected Return of Stock A =0.1*10%+0.2*2%+0.4*12%+0.2*20%+0.1*38% =14%
Expected Return of Stock B =0.1*35%+0.2*0%+0.4*20%+0.2*25%+0.1*45% =21%

Option A is the correct option